please "simplify" this problem.

3m^6 over 6m^6

thanks

Answer by RAY100(1637)

(3m^6) / (6m^6)
remember to divide powers with common bases, we subtract exponents
,
m^6/m^6 = m^(6-6) =m^0 =1 (anything raised to zero power is 1)
3/6 = 1/2
,
answer is,,, 1/2